OpenStudy (anonymous):

in a bag filled with only red, green and blue marbles, 1/4 of the marbles in a bag are red and 2/3 of the remaining marbles are green. if there are t marbles in total,in terms of t what percent of the marbles are blue ??

OpenStudy (anonymous):

(1/4)t + (2/3)t + (x)t = t. 1/4 + 2/3 + x = 1. Solve for x. Answer will be xt.

OpenStudy (anonymous):

Actually scratch that. Just saw "2/3 of the REMAINING" are green. New equation needed.

OpenStudy (anonymous):

so whats the new equation?@tcarroll010

OpenStudy (anonymous):

(1/4)t + (2/3)(3/4)t + (x)t = t. 1/4 + (2/3)(3/4) + x = 1. Solve for x. Answer will be xt. I added the additional factor of 3/4 in the second term.

OpenStudy (anonymous):

lets say there are 24 marbles in the bag. one quarter of them or 6 are red leaving 18 two thirds of the remaining 18 are green and 2/3 of 18 is 12 that leaves 6 unaccounted for marbles. they must be blue so it is 6 out of 24 or 1/4 or 25%